OpenStudy (anonymous):

Every month, a girl gets allowance. Assume last year she had no money, and kept it up to now. Then she spends 1/2 of her money on clothes, then 1/3 of the remaining money on games, and then 1/4 of the remaining money on toys. After she bought all of that, she had $7777 left. Assuming she only gets money by allowance, how much money does she earn every month?

OpenStudy (anonymous):

Work backwards... 7777 * 4/1 = 31108 * 3/1= 93324 * 2 = 186648 /12 = 15554